### 05 Feb 2024 v1

# A Survey on Transformer Compression

Abstract
Large models based on the Transformer architecture play increasingly vital roles in artificial intelligence, particularly within the realms of natural language processing (NLP) and computer vision (CV). Model compression methods reduce their memory and computational cost, which is a necessary step to implement the transformer models on practical devices. Given the unique architecture of transformer, featuring alternative attention and Feedforward Neural Network (FFN) modules, specific compression techniques are required. The efficiency of these compression methods is also paramount, as it is usually impractical to retrain large models on the entire training dataset. This survey provides a comprehensive review of recent compression methods, with a specific focus on their application to transformer models. The compression methods are primarily categorized into pruning, quantization, knowledge distillation, and efficient architecture design. In each category, we discuss compression methods for both CV and NLP tasks, highlighting common underlying principles. At last, we delve into the relation between various compression methods, and discuss the further directions in this domain.
Index Terms:
Model Compression, Transformer, Large Models, Computer Vision, Natural Language Processing
1 Introduction
Deep neural networks have become indispensable in numerous artificial intelligence applications, whose architectures encompass diverse formulations, such as multilayer perceptron (MLP), convolutional neural network (CNN), recurrent neural network (RNN), long short-term memory (LSTM), Transformers, etc. In recent times, transformer-based models have emerged as the prevailing choice across various domains, including both natural language processing (NLP) and computer vision (CV) domains. Considering their strong scaling ability, most of the large models with over billions of parameters are based on the transformer architecture, which are considered as foundational elements for general artificial intelligence (AGI) [1, 2, 3, 4, 5, 6].
While large models have demonstrated significant capabilities, their exceptionally vast sizes pose challenges for practical development. For instance, the GPT-3 model has 175 billion parameters and demands approximately about 350GB memory model storage (float16). The sheer volume of parameters and the associated computational expenses necessitate devices with exceedingly high memory and computational capabilities. Directly deploying such models will incur substantial resource costs and contributes significantly to carbon dioxide emissions. Moreover, on edge devices like mobile phones, the development of these models becomes impractical due to their limited storage and computing resource.
Figure 1: Transformer-based models have emerged as the predominant architectures in both natural language processing (NLP) and computer vision (CV) domains, garnering a surge in publications. As these models tend to possess substantial dimensions, the imperative arises to compress their parameters and streamline computational redundancies. This compression is essential for facilitating efficient implementation on practical platforms, ensuring the feasibility of deploying Transformer models in real-world applications.
Model compression is an effective strategy for mitigating the development costs associated with transformer models. This approach, grounded in the principle of reducing redundancy, encompasses various categories, including pruning, quantization, knowledge distillation, efficient architecture design, etc. Network pruning directly removes redundant components, such as blocks, attention heads, FFN layers or individual parameters. Diverse sub-models can be derived by employing different pruning granularity and pruning criteria. Quantization reduces the development cost by representing model weights and intermediate features with lower bits. For example, When quantizing a full-precision model (float32) into 8-bit integers, the memory cost can be reduced by a factor of four. According the computational process, it can be divided into post-training quantization(PTQ) or quantization-aware training (QAT), in which the former only incurs limited training costs and more efficient for large models. Knowledge distillation serves as a training strategy, which transfers knowledge from a large model (teacher) to a smaller model (student). The student mimics the behavior of the teacher by emulating the model’s output and intermediate features. Notably, for advanced models like GPT-4, accessible only through APIs, their generated instructions and explanations can also guide the learning of the student model [7, 8].In addition to obtaining models from predefined large models, some methods yield efficient architectures by directly reducing the computational complexity of attention modules or FFN modules. Combining different methods enables extreme compression. For instance, Han et al. [9] combined network pruning, quantization, and Huffman coding, achieving an impressive 49$\times$ compression rate on a conventional VGGNet [10].
Regarding transformer models, their compression strategies exhibit distinct characteristics. Different from other architectures like CNN or RNN, the transformer features a unique design with alternative attention and FFN modules. The former captures the global information by calculating the attention map over different tokens while the other extracts information from each token respectively. This specific architecture can inspire tailored compression strategy for optimal compression rates. Besides, the efficiency of compression method becomes especially important for such large models. Due to the high computational cost of large model, it is usually unaffordable to retrain the whole model on the original training set. Some training-efficient methods like post-training compression are more preferable.
In this survey, we aim to comprehensively investigate how to compress these transformer models (Figure 1), and the methods are categorized by quantization, knowledge distillation, pruning, efficient architecture design, etc. In each category, we investigate the compression methods for NLP and CV domains, respectively. Table I summarizes the main compression categories and lists representative method suitable for large transformer models. Though NLP and CV are usually treated as very different domains, we observe that their models compression methods actually share the similar principles. Finally, we discuss the relationship between different compression methods and post some future research directions.
The rest of the paper is organized as follows. Section 2 introduces the fundamental concept of transformers. Following this, Section 3 provides an in-depth discussion on compression methods that preserve the architecture, encompassing quantization and knowledge distillation—techniques that maintain the model’s architecture. Section 4 delves into further exploration of architecture-preserving compression, including pruning and efficient architecture design. Additional transformer compression methods are explored in Section 5. Finally, Section 6 concludes the compression method and discuss future research directions.
2 Concept of Transformer
The transformer architecture is firstly proposed to tackle tasks like machine translation [11]. A standard Transformer architecture contains main blocks, multi-head attention (MHA) and feed-forward networks (FFN). The attention is formulated as
$\mathrm{Attention}(Q,K,V)=\mathrm{softmax}(\frac{QK^{T}}{\sqrt{d}})V,$ (1)
where $Q$, $K$, $V$ are query, key and value matrix, respectively. $d$ is the feature’s dimension. The multi-head attention jointly extracts information from diverse subspaces, which is the concatenation of different heads, i.e.,
$\mathrm{MultiHead}(Q,K,V)=\mathrm{Concat}(\mathrm{head_{1}},...,\mathrm{head_{h}})W^{O},$ (2)
$\text{where}~{}\mathrm{head_{i}}=\mathrm{Attention}(QW^{Q}_{i},KW^{K}_{i},VW^{V}_{i}).$
$W^{Q}$, $W^{K}$, $W^{V}$, $W^{O}$ are the corresponding parameter matrices. The FFN module transforms features from each token independently. It is usually constructed by stacking two FC layers with activation functions, i.e.,
$\mathrm{FFN}(x)=\phi(xW_{1}+b_{1})W_{2}+b_{2},$ (3)
where $x$ is input feature, $\phi$ is activation function (e.g., GeLU). $W_{1}$, $W_{2}$, $b_{1}$, $b_{2}$ are the weight and bias parameters in FC layers. The MHA and FFN module are stacked alternatively to construct the whole model.
The Transformer architecture has strong scaling ability to construct extremely large models with several billion or trillion parameters. It supports most of the prevalent large models in NLP, CV and multiple modality domains. For example, the well-known large language models (e.g., GPT-series [4, 2], LLAMA [1],Pangu [5, 6]) are its decoder-only variants. By simply splitting an image into multiple patches, it can be used to tackle vision tasks [12, 13, 14]. The multiple model like CLIP [15], BLIP [16], LLaVA [17] also use transformer as the backbones.
TABLE I: Representative compression method for transformer models.
Category Sub-category Method Highlights Publication
Quantization NLP SmoothQuant [18] Training-free, Smooth outliers, Equivalent transformation ICML 2023
OmniQuant [19] Weight clipping, learnable transformation, block-wise Arxiv 2023
QLoRA [20] Parameter-efficient fine-tuning, memory management Arxiv 2023
CV PTQ-ViT [21] Self-attention preservation, mixed-precision NeurIPS 2021
FQ-ViT [22] Fully-quantized, log2 quantization, power-of-two factor IJCAI 2022
OFQ [23] Confidence-guided annealing, query-key reparameterization ICML 2023
Knowledge Distillation NLP DistilBERT [24] Small version of BERT, trained with logits of the teacher NeurIPS 2019
MiniLM [25] Mimicking attention distribution and value-relation of teacher NeurIPS 2020
Lion [7] Adversarial distillation: imitation, discrimination, generation EMNLP 2023
CV DeiT [26] Hard labels, novel distillation token in ViTs ICML 2021
TinyViT [27] Large-scale pretraining data, encoded data augmentation ECCV 2022
ManifoldKD [28] Patch-level, batch-level manifold information NeurIPS 2022
Pruning NLP LLM Pruner [29] Structured, Coupled-module identification NeurIPS 2023
Sheared LLaMA [30] Structured, Pre-defined tagert, dyanmic data loading NeurIPS 2023
Dynamic Context Pruning [31] sigmoid-based context selection, KV-cache aware NeurIPS 2023
CV ViT-Slim [32] Structured, Single-shot architecture search CVPR 2022
Patch Sliming [33] Top-down unimportant patch removing CVPR 2022
X-pruner [34] Structured, Class-aware, Layer-wise fully differentiable pruning CVPR 2023
Efficient Architecture NLP PaLM [35] SwiGLU Activation in FFN, densely activated JMLR 2023
RetNet [36] Training parallelism, low-cost inference, parallel, recurrent Arxiv 2023
Reformer [37] Efficient attention, locality-sensitive hashing ICLR 2020
CV Swin [38] Hierarchical structures, shifted local window attention ICCV 2021
MetaFormer [39] Non-parametric pooling as basic token mixing CVPR 2022
MLP-Mixer [40] Architecture based exclusively on multi-layer perceptrons NeurIPS 2021
3 Architecture Preserved Compression
3.1 Quantization
3.1.1 Overview of Quantization
Quantization is a crucial step for deploying transformers on various devices, especially GPUs and NPUs, that design specialized circuits for low-precision arithmetic. During the quantization process as Equation 4, a floating-point tensor $x$ is converted to the integer one $x_{int}$ with corresponding quantization parameters (scale factor $s$ and zero point $z$), then the integer tensor $x_{int}$ could be quantized back to floating-point $x_{quant}$ but causes some precision error compared with original $x$, i.e.,
$x_{int}$ $=\textrm{Clamp}(\lfloor x/s\rceil+z,0,2^{b}-1),$ (4)
$x_{quant}$ $=s(x_{int}-z),$
[…]
$y$ $=\textrm{MatMul}(e,\textrm{w})\approx\textrm{MatMul}(e_{quant},\textrm{w}_{quant})$ (5)
$=\textrm{MatMul}(s_{e}(e_{int}-z_{e}),s_{w}\textrm{w}_{int})$
$=s_{e}s_{w}\textrm{MatMul}(e_{int},\textrm{w}_{int})+C,$
where $s_{w},s_{e},z_{e}$ are quantization parameters of weights and input embedding, and $e_{int}$ and $\textrm{w}_{int}$ are the integer input and weights, which calculated by Equation 4. $C$ could be pre-computed with $s_{e},z_{e},s_{e}$ and $\textrm{w}_{int}$. Thus the floating-point multiplication could be accelerate with efficient integer multiplication in the inference. To minimize the performance degradation of quantized models, different optimization methods have been proposed and they can be divided into two categories: (1) Post-training quantization (PTQ) [21, 41, 22, 42, 43, 44, 45] , mainly focuses on optimizing the quantization parameters of weights and activations with a few unlabeled calibration data, some latest methods also explore adaptive rounding for weight quantization. (2) Quantization-aware training (QAT) [46, 47, 48, 49, 50, 51, 23, 52, 53, 54, 55, 56], inserts the quantization nodes into networks and conducts training with complete training data, where all the weights and quantization parameters are optimized together. In this section, we systematically introduce the research of model quantization on transformer-based vision models and large language models, as shown in Figure 2.
Figure 2: The overview of quantization for transformers. The top demonstrates the different problems that addressed in existing works for computer vision and natural language processing, and the bottom shows a normal INT8 inference process of a standard transformer block.
TABLE II: Comparison of different PTQ and QAT methods for transformer-based vision models. W/A denotes the bit-width of weight and activation and the results show the top-1 accuracy on ImageNet-1k validation set. * represents for mixed-precision.
Method of PTQ W/A (bit) ViT-B DeiT-S DeiT-B Swin-S Swin-B
Full Precision 32/32 84.54 79.85 81.80 83.23 85.27
PTQ-ViT [21] 8*/8* 76.98 78.09 81.29 - -
PTQ4ViT [41] 8/8 84.54 79.47 81.48 83.10 85.14
FQ-ViT [22] 8/8 83.31 79.17 81.20 82.71 82.97
APQ-ViT [42] 8/8 84.26 79.78 81.72 83.16 85.16
NoiseQuant [44] 8/8 84.22 79.51 81.45 83.13 85.20
PTQ-ViT [21] 6*/6* 75.26 75.10 77.47 - -
PTQ4ViT [41] 6/6 81.65 76.28 80.25 82.38 84.01
APQ-ViT [42] 6/6 82.21 77.76 80.42 84.18 85.60
NoiseQuant [44] 6/6 82.32 77.43 80.70 82.86 84.68
RepQ-ViT[45] 6/6 83.62 78.90 81.27 82.79 84.57
APQ-ViT [42] 4/8 72.63 77.14 79.55 80.56 81.94
PTQ-ViT [21] 4*/4* - - 75.94 - -
APQ-ViT [42] 4/4 41.41 43.55 67.48 83.16 85.16
RepQ-ViT[45] 4/4 68.48 69.03 75.61 79.45 78.32
Method of QAT W/A (bit) DeiT-T DeiT-S DeiT-B Swin-T Swin-S
Full Precision 32/32 72.20 79.85 81.80 81.20 83.23
I-ViT [50] 8/8 72.24 80.12 81.74 81.50 83.01
Q-ViT [48] 4/4 72.79 80.11 - 80.59 -
AFQ-ViT [48] 4/4 - 80.90 83.00 82.50 84.40
Quantformer [46] 4/4 69.90 78.20 79.70 78.30 81.00
OFQ [23] 4/4 75.46 81.10 - 81.88 -
VVTQ [56] 4/4 74.71 - - 82.42 -
Q-ViT [48] 3/3 69.62 78.08 - 79.45 -
AFQ-ViT [48] 3/3 - 79.00 81.00 80.90 82.70
Quantformer [46] 3/3 65.20 75.40 78.30 77.40 79.20
OFQ [23] 3/3 72.72 79.57 - 81.09 -
AFQ-ViT [48] 2/2 - 72.10 74.20 74.70 76.90
Quantformer [46] 2/2 60.70 65.20 73.80 74.20 76.60
OFQ [23] 2/2 64.33 75.72 - 78.52 -
(a) ViT-B_16
(b) ViT-B_16-224
(c) ViT-L_16
(d) ViT-L_16-224
(e) OPT-13B
(f) OPT-30B
(g) OPT-66B
(a) ViT-B_16
(b) ViT-B_16-224
(c) ViT-L_16
(d) ViT-L_16-224
(e) OPT-13B
(f) OPT-30B
(g) OPT-66B
(h) OPT-175B
Figure 3: Inference latency of the ViT and OPT using FasterTransformer on NVIDIA A100-80GB GPUs. The data of OPT is from [18].
3.1.2 Quantization for Transformer-Based Large Language Models
Before 2023, the study of quantization [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] on transformer-based NLP almost focuses on BERT architecture. With the popularity of pretrained large language models [68, 1, 2], researchers [69, 70, 71] explore how to quantize the transformer with Billion-level parameters and explores more efficient quantization schemes with limited data and computational overhead.
Post-training Quantization. Based on the analysis of outliers for quantized transformers, Outlier Suppression [72] proposes to migrate the gamma of LayerNorm to the next module for that the gamma amplifies the outliers in the output and causes large quantization error, and then clips the tokens in a coarse-to-fine procedure. MREM [59] focuses on reducing the computational cost of quantization, presents to minimize the output quantization error for all modules in parallel manner. Zeroquant [71] finds that the performance degradation comes from the different dynamic range of tokens and weights, proposes to adopt group-wise quantization for weights and token-wise quantization for activations, and utilizes knowledge distill layer by layer. For large transformers with billion-level parameters, the outlier is still the main reason that resulting large accuracy degradation for quantized models. To address that, LLM.int8() [69] represents the activations and outliers of weights with 16-bit and performs 8-bit vector-wise quantization for weight tensor, but the acceleration is limited and even worse due to the irregular quantization granularity [18]. GPTQ [73] also only quantizes weight parameters as LLM.int8(), but adopts unify quantization strategy. GPTQ quantizes the weights based on approximate second-order information, which could get much accurate quantized weight in a few hours. AWQ [74] proposes to search the optimal per-channel scale factors by observing the distribution of activations instead of weights, thus the LLMs could remain the generalist abilities for different domains and modalities. Outlier Suppression+ [75] explores more accurate outliers suppression scheme with channel-wise shifting and scale, which helps align the range of different activation channels and scale down the outliers. The shifting and scale factors could also be merged with other weight parameters. Similarly, SmoothQuant [18] and QLLM [76] propose the mathematically equivalent per-channel scaling transformation that migrates the quantization difficulty from activations to weights. To further improve the performance of quantized LLMs, QLLM also learns a low-rank parameters by minimizing the reconstruction error of the outputs between floating-point and quantized LLMs with limited calibration data. Similarly, RPTQ [77] utilizes a reorder-based scheme that rearranging the channels of activations with similar range and quantizes them in cluster, and then migrates the scale into LayerNorm and weights of linear layers without extra computational overhead in the inference. Based on scale migration, OmniQuant [19] further proposes a learnable PTQ methods module by module, where the weight clipping parameters and transformation scale of activations are optimized with gradient decent. Like GPTQ, SqueezeLLM [78] also focuses on quantizing weights for the memory bandwidth and proposes to search the optimal bit based on second-order information. Besides, SqueezeLLM does not suppression the outliers and sensitive weight values but store them in a efficient sparse format for more accurate quantized LLMs. Different from the previous methods, SignRound [79] proposes to optimize quantized LLMs from the perspective of adaptive rounding. SignRound designs block-wise tuning using signed gradient descent to learn the weight rounding, which greatly help the output reconstruction of each block.
TABLE III: Perplexity (PPL) comparison of different quantization methods on WikiText2. W/E denotes weights and embedding, respectively. * represents adopting different methods to get the perplexity, please refer to the corresponding papers.
Method PTQ Bits LLaMA
(W-E) 7B 13B 30B 65B
FP16 $\times$ 16/16 5.68 5.09 4.10 3.56
SmoothQuant* [18] 8/8 11.56 10.08 7.56 6.20
LLM-QAT* [80] $\times$ 8/8 10.30 9.50 7.10 -
OS+ [75] 6/6 5.76 5.22 4.30 3.65
OmniQuant [19] 6/6 5.96 5.28 4.38 3.75
QLLM [76] 6/6 5.89 5.28 4.30 3.73
SqueezeLLM [78] 4/16 5.79 5.18 4.22 3.76
SignRound [79] 4/16 6.12 5.32 4.52 3.90
OmniQuant [19] 4/16 5.86 5.21 4.25 3.71
LLM-QAT* [80] $\times$ 4/8 10.90 10.00 7.50 -
OS+ [75] 4/4- 14.17 18.95 22.61 9.33
OmniQuant [19] 4/4 11.26 10.87 10.33 9.17
QLLM [76] 4/4 9.65 8.41 8.37 6.87
PEQA [81] $\times$ 4/4 5.84 5.30 4.36 4.02
SqueezeLLM [78] 3/16 6.32 5.60 4.66 4.05
OmniQuant [19] 3/16 6.49 5.68 4.74 4.04
PEQA [81] $\times$ 3/3 6.19 5.54 4.58 4.27
OmniQuant [19] 2/16 15.47 13.21 8.71 7.58
Quantization-aware training. Q-BERT [57] is the early work that conducts quantization-aware training on transformer-based architecture for natural language processing. Inspired by HAWQ [82], Q-BERT searches mixed-quantization setting based on the second order Hessian information and adopts group-wise quantization to partition parameter matrix into multiply groups to reduce accuracy degradation. I-BERT [62] not only quantizes the linear and self-attention layer, but also designs an integer-only inference scheme for the nonlinear operations (GELU, SoftMax and LayerNorm) in transformers, which also inspires the FQ-ViT and I-ViT for the quantization of vision transformer. Specifically, I-BERT proposes to use the second-order polynomial approximate for GELU and exponential function of SoftMax, and calculate the standard deviation of LayerNorm based on Newton’s Method, thus I-BERT could achieve faster inference compared to the baseline and normal quantization methods. To compensate for the disadvantages of the hand-crafted heuristics in Q-BERT, AQ-BERT [66] proposes an automatic mixed-precision quantization scheme to learn the bit-width and parameters of each layer simultaneously, which is inspired from differentiable network architecture searching. AQ-BERT achieves better results than Q-BERT and more suitable for resource-limited devices. Except the QAT schemes for BERT, there are only a few papers that conducts quantization-aware training on large language models for the huge training overhead. PEQA [81] and QLoRA [20] explore parameter-efficient fine-tuning techniques to train the quantized LLMs, the former only optimizes the scale factors while freezing the quantized weights, and the later adopts 4-bit NormalFloat, double quantization and a paged optimizers for normally distributed weights and reduce memory footprint during training. LLM-QAT [80] proposes a data-free quantization-aware training methods, where the training data are generated from the original pretrained large language model. LLM-QAT not only quantizes all the linear layers and self-attention, but also quantizes the KV cache with cross-entropy based logits distillation.
3.1.3 Quantization for Transformer-Based Vision Models
[…]
Figure 4: The taxonomy of knowledge distillation used for large transformer-based models.
Quantization-aware training. When compressing vision transformers to extremely low-bit precision, PTQ can not optimize the large quantization error with limited calibration images and suffers from significant performance reduction, thus QAT are urgently required for more accurate low-bit vision transformers. Q-ViT [48] finds that MSA and GELU are much sensitive to quantization, thus proposes a fully differentiable quantization method, which adopts head-wise bit-width and switch scale during quantization searching process. Quantformer [46] takes the self-attention rank into consideration, and proposes to maintain the consistency between quantized and full-precision vision transformers. Besides, Quantformer presents the group-wise strategy to quantize feature of patches in different dimensions, where each group adopts different quantization parameters and the extra computation overhead is negligible. Based on the observation that the sever performance degradation suffers from the quantized attention map, AFQ-ViT [47] designs an information rectification module and a distribution guided distillation during quantization training. The former one helps recover the distribution of attention maps with information entropy maximization in the inference, and the later one reduces the distribution variation with attention similarity loss in the backward. Similar with FQ-ViT [22], I-ViT [50] also explores the integer-only quantization scheme for ViTs. I-ViT designs Shiftmax, ShiftGELU and I-LayerNorm to replace the vanilla modules with bit-wise shift and integer matrix operations in the inference, achieving 3.724.11×3.72-4.11$\times$3.72 - 4.11 × acceleration compared to the floating-point model. OFQ [23] finds that weight oscillation causes unstable quantization-aware training and leads to sub-optimal results, and the oscillation comes from the learnable scale factor and quantized query and key in self-attention. To address that, OFQ proposes statistical weight quantization to improve quantization robustness, and freezes the weights with high confidence and calms the oscillating weights with confidence-guided annealing. For the query and key in the self-attention, OFQ presents query-key reparameterization to decouple the negative mutual-influence between quantized query and key weights oscillation. Similarly, VVTQ [56] analyzes ViT quantization from the perspective of variation, which indicates that the data variance within mini-batch are much harmful to quantization and slow down the training convergence. To reduce the impact of variations in the quantization-aware training, VVTQ proposes a multi-crop knowledge distillation-based quantization methodology, and introduces module-dependent quantization and oscillation-aware regularation to enhance the optimization process.
We summarize the results of these PTQ and QAT methods in Table II, most schemes conduct PTQ on 8bit and 6-bit, there are server degradation when quantized to 4-bit. With the complete training data, QAT could push the ViT quantization to 4-bit and less, some methods even shows better results than floating point models, which proves that quantization-aware training could tap into the true accuracy of quantized models efficiently. And Figure 3 shows the latency results of ViT models using FasterTransformer [^1] https://github.com/NVIDIA/FasterTransformer.
3.1.4 Discussion
With the presence of extreme distributions and outliers, quantization for transformers is much more difficult than that for convolutional neural networks. To recover the performance of quantized models, various methods are proposed to address the quantization-unfriendly components in transformers. Specifically, For vision tasks, the existing methods optimizes the quantized transformers in three ways: remaining the self-attention rank, rectifying the distribution of extreme distribution and addressing the weight oscillation and data variation in quantization-aware training, as shown in Figure 2. And for natural language processing task, most schemes aims to process the outlier of weights and activations. Besides, the training overhead is usually unacceptable for large language models with billion-level parameters, thus quantizing module by module and conducting parameter-efficient fine-tune are much popular. However, when compressing to extremely low bit-width, the quantized transformers suffer significant performance degradation, which are far away from the floating-point models, thus how to build a more accurate and low-bit transformers is still a hard problem to be solved.
3.2 Knowledge Distillation
In this section, we will introduce knowledge distillation frameworks used in compressing transformer based foundation models, including both language and vision models.
3.2.1 Overview of Knowledge Distillation
Knowledge distillation (KD) aims to train student networks by compressing [83, 84, 85] or transferring [87, 88, 86] knowledge from teacher networks. In this paper, our main focus lies on distillation methods proposed to achieve a compact student model while preserving satisfactory performance compared to a cumbersome teacher model. The student models typically have narrower and shallower architectures, making them more suitable for deployment on resource-limited systems.
We will first discuss two different kinds of knowledge distillation in the following: logits-based methods [87] which convey knowledge on the logits level and hint-based methods [89] which convey knowledge through intermediate features. To illustrate the objective of logits-based KD in classification tasks, we denote the logits output of the teacher and student network as $z^{\bf s},z^{\bf t}\in\mathbb{R}^{C}$, where $C$ represents the number of classes. Neural networks often generate class probabilities by applying a softmax function to the logits, converting them into probabilities $p^{\bf s},p^{\bf t}$ as follows:
$p_{i}^{\bf s}=\frac{\exp(z_{i}^{\bf s})}{\sum_{j=1}^{C}\exp(z_{j}^{\bf s})},\quad p_{i}^{\bf t}=\frac{\exp(z_{i}^{\bf t})}{\sum_{j=1}^{C}\exp(z_{i}^{\bf t})},$ (6)
With above probabilities, the logits-based KD minimizes the KL divergence between the probabilities of the teacher and student model as:
$L_{logits}=KL(p^{\bf t}||p^{\bf s})=\sum_{j=1}^{C}p_{j}^{\bf t}log(\frac{p_{j}^{\bf t}}{p_{j}^{\bf s}}),$ (7)
As for hint-based KD, given the intermediate features of the teacher and student network, $F^{\bf s},F^{\bf t}\in\mathbb{R}^{H\times W\times C}$, the corresponding loss is formulated as:
$L_{hint}=\mathcal{H}(F^{\bf s},F^{\bf t})=||F^{\bf t}-\phi(F^{\bf s})||^{2},$ (8)
where $\phi$ is a function used to ensure that the student features have the same shape as the teacher features. $\mathcal{H}$ represents the chosen metric function, here we provide an example using mean squared error.
In addition to the two mainstream KD methods mentioned above, we will also discuss an API-based method, where only the teacher’s generated outputs are accessible in today’s large language models (LLMs).
3.2.2 KD for Transformer-Based Large Language Models
Logits-based KD. Tang et al. [90] employed knowledge distillation to compress BERT[91], a large language model, into a much lighter bidirectional long short-term memory network (BiLSTM) [92] for natural language processing (NLP) tasks. The distillation process aimed to minimize the mean squared error (MSE) loss between the student’s logits and the teacher’s logits. Following distillation, the shallow BiLSTM-based model achieved comparable results to ELMo[93], but with approximately 100$\times$ fewer parameters and a 15$\times$ faster inference speed. Similarly, DistillBERT [24] initialized the shallower student with teacher’s parameters, and minimized the soft target probabilities between the teacher and the student, also known as word-level KD. SeqKD [94] fine-tuned the student model on the sequence-level teacher-generated data. MixKD [95] extended the idea of encouraging the student to mimic the teacher’s logits to the linear interpolation of example pairs [96]. Theoretical analysis has shown that, under reasonable conditions, MixKD can effectively reduce the gap between generalization error and empirical error. Turc et al. [97] demonstrated the ongoing significance of pre-training, even when employing smaller architectures. They introduced a methodology called Pre-trained Distillation (PD), which begins with pre-trained compact student models. Subsequently, it explores the transfer of task knowledge from large fine-tuned models using conventional logits-based KD. In contrast, MINILLM [98] highlighted a limitation of conventional logits-based KD that minimize forward Kullback-Leibler divergence (KLD) in free-run generation. They found that such approaches can lead the student model to overestimate the low-probability regions of the teacher’s distribution. To address this issue, MINILLM proposed replacing the forward KLD objective (as shown in Eq. 7) with reverse KLD, i.e., $reverse~{}\mathrm{KLD}:=KL(p^{\bf s}||p^{\bf t})$, which is better suited for knowledge distillation on generative language models. GKD [99] identified a challenge in distilling knowledge into auto-regressive student models, which is the train-test distribution mismatch. Specifically, partial sequences encountered by the student during generation phase can be significantly different from the ones observed during the training phase. To address this issue, GKD trained the student on its self-generated output sequences by leveraging feedback (logits) from the teacher on such sequences. It also facilitated the integration of distillation with RL fine-tuning of language models. Huang et al. [100] further combined the hint-based KD with multitask in-context learning.
Hint-based KD. Sun et al. [101] proposed patient knowledge distillation to help shallow BERT student learn from multiple intermediate features [102, 103, 104, 25] of the deep BERT teacher. Li et al. [105] utilized hints extracted from both intermediate hidden states and attention distributions to enhance the training of Non-AutoRegressive Translation (NART) models. The distilled NART models attained performance similar to a powerful LSTM-based AutoRegressive Translation (ART) baseline in various machine translation tasks. Notably, the distilled NART models exhibited a speed improvement of 17$\times$ over their ART counterparts. Similarly, Mukherjee et al. [106] employed MBERT [107] as a teacher model to guide the training of smaller BiLSTM models, resulting in a significant 51$\times$ speed improvement for batch inference. MobileBERT [103] implemented two objective functions to distill knowledge from a BERT teacher incorporated with inverted bottleneck, including attention distributions and hidden states, to a slimmed-down version of BERT as the student model. In addition to these objectives, TinyBERT [104] extended the distillation process by also transferring knowledge from the teacher’s logits to the student model. Wang et al. [25] proposed MINILM to train the student networks to mimic both the attention distributions and the scaled dot-product between values of teacher models. It retains more than 99% accuracy on GLUE benchmark tasks using 50% of the Transformer parameters and computations of the teacher model.
API-based KD. Today’s language models have reached unprecedented scale [4, 108, 109] in terms of computation and model parameters, as evident from the significant advancements made in large language models (LLMs) like GPT-4 [2]. However, the power of such models is only accessible through their APIs, which limits the ability to utilize conventional KD methods that rely on intermediate features or logits of the teacher model. In this section, we will discuss API-based KD, also known as black-box KD, where only the teacher’s final generations are accessible for the purpose of distillation.
Several recent studies [110, 111, 112, 113, 114, 7, 115, 8, 116] have showed the promising results in fine-tuning small models on the outputs generated by LLMs. Specifically, they first prompted a very large teacher model, such as GPT-3 with 175B parameters, to solve complex questions via chain-of-thought (CoT) reasoning [117]. Then, they used the generated instructions [7, 115], explanations [8] or rationales [110] to fine-tune a much smaller student. Furthermore, Zhu et al. [118] proposed PaD, which utilizes program-aided reasoning, such as an additional Python interpreter, to help student models overcome faulty steps in the CoT reasoning with automated error checking. Shridhar et al. [119] introduced Socratic CoT that trains a combination of two small distilled models: a problem decomposer and a subproblem solver. These models work together to decompose and solve complex problems effectively. Fu et al. [120] demonstrated that specializing the small student model’s abilities for the specific reasoning task, by transferring the knowledge in larger teacher model’s generic directions, yields promising results. Jiang et al. [7] identified a shortcoming in the aforementioned approaches, as they lacked the ability to incorporate feedback for identifying challenging instructions where the student model’s performance falls short. To address this limitation, they proposed a three-stage adversarial loop of distillation that incorporates feedback and addresses these challenging instructions.
Other KDs. Xu et al. [121] compressed the BERT model by progressive module replacing. The original model was first divided into multiple modules, and more compact substitutes are created for these modules. Then, they randomly substituted the original modules with their corresponding substitutes and train the compact modules without any additional loss functions.
Jha et al. [122] addressed the challenge of fitting both the student and teacher models, which consist of hundreds of parameters, into GPU memory, making traditional knowledge distillation infeasible. To overcome this limitation, they proposed a teacher-free task-agnostic distillation method which uses a truncated version (student) of the large model (teacher) for initialization [123] and then continuing the pre-training of student, without employing any distillation loss.
TABLE IV: Comparison with previous transformer-based language model distillation approaches. The GLUE score is averaged on 8 tasks, i.e., SQuAD2, MNLI-m, SST-2, QNLI, CoLA, RTE, MRPC, QQP.
  Models Distillation Teacher ##\## Layer ##\## Params Speed Up GLUE
[…]
  • ${}^{\dagger}$ The corresponding data is from [124].
  • ${}^{\ddagger}$ The corresponding data is from [98]. They reported F1 for SQuAD 2.0, and accuracy for other datasets.
3.2.3 KD for Transformer-Based Vision Models
Logits-based KD. While conventional KD in vision tasks typically involves using soft labels [87, 125, 84] generated by a powerful teacher network, Touvron et al. [126] made a noteworthy discovery. Their experiments on the ImageNet benchmark revealed that employing hard labels, which are essentially the maximum of the score values predicted by the teacher model, led to superior results. Additionally, they introduced a novel distillation token into vision transformers (ViTs)[127]. TinyViT et al.[27] observed that smaller ViTs could benefit from larger teacher using massive pre-training data, i.e., pre-training (distilling) student on ImageNet-21k while fine-tuning student on ImageNet-1k. And they proposed TinyViT to save computational memory by storing data augmentation information and logits in advance for large teacher models. Ren et al. [128] argued that student accuracy is primarily influenced by the inductive bias of the teacher models rather than their accuracy. To address this, they introduced a cross-inductive bias distillation approach, which involves distilling the student model with multiple teachers exhibiting distinct architectural inductive biases. To enhance the utilization of the inductive bias in CNNs and offer more image-friendly guidance, CSKD [129] removed the pooling operation after the final feature map and treated the features at each position as individual samples. Subsequently, these position-based features were input into the classifier, leading to the generation of dense predictions for each corresponding position. The hard labels derived from CNN’s dense predictions were used as target labels, and cross-entropy served as the loss function for spatial knowledge transfer.
Hint-based KD. Chen et al. [130] introduced a two-stage DearKD framework. In the initial stage, knowledge distillation is employed from the intermediate layers of the CNN to help transformer-based student capture the inductive biases. Subsequently, the second stage trains the student without distillation. Hao et al. [28] proposed a fine-grained manifold distillation method to fully utilize the patch-level and batch-level information inside transformer-based architectures. Specifically, they manually selected intermediate layers from teacher-student pair and distilled the student via decoupled relation maps. ViTKD et al. [131] delved into the characteristics of feature maps in ViTs and formulated three practical guidelines for hint-based distillation in ViTs, namely: (i) generating hints is more effective than mimicking in deeper layers, (ii) distillation on the shallower layers is also suitable for ViTs with mimicking, and (iii) features from the feed-forward network (FFN) are better than those from the multi-head attention (MHA) for the distillation process.
3.2.4 Discussion
Transformer-based language models have shown strong performance on several NLP tasks including text classification, machine translation, speech recognition, question-answering, dialogue systems, etc. Knowledge distillation is a common and practical technique used to reduce the high computational resource demand of LLMs. Different models, such as T5 [132], BERT [91], LLAMA [1], GPT-3 [4], and GPT-4 [2], exhibit varying performance on different NLP downstream tasks. Therefore, selecting appropriate teachers and methods for model compression is crucial. For example, when compressing the BERT series, we compare various distillation methods in Table IV. In contrast to using only logits, hint-based KD methods can transmit richer intermediate layer information to the student, making the learning process easier for the student and yielding better results. However, performing layer-to-layer knowledge distillation sometimes requires carefully designed layer mappings between the teacher and student models. In certain domains where generative models excel, such as reasoning and language understanding, the effectiveness of strong teacher models can only be accessed through their APIs. The exploration of guiding LLMs to generate better outputs for distillation is still in its early stages.
One of the strengths of ViT models is their ability to scale to high parametric complexity, but this requires substantial computational resources and comes at a high cost. KD can be used to transfer knowledge into more compact student models, but there are still challenges in the context of vision that require further research. The first challenge is related to training costs. Both logits-based and hint-based KD methods require large GPU memory for the distillation process. Vision tasks differ from NLP as the input for vision is in the form of images, which are significantly larger in size compared to the limited sequence length used in NLP tasks. During distillation, even though the teacher network doesn’t require backward propagation, each forward pass consumes a significant amount of GPU memory due to the activations of intermediate features. Therefore, finding ways to reduce training costs and shorten training times is an area that needs exploration. Additionally, choosing the most suitable teacher model for transformer-based students is an open question. For instance, in the context of ViTs, determining whether a CNN or a transformer teacher is more effective for the student model is a potential research avenue. Moreover, as proposed by [84], there is a necessity for designing and evaluating KD approaches within practical scenarios, moving away from the limitations of small-scale datasets. This ensures that KD methods are applicable and effective in real-world, large-scale settings.
4 Architecture Adaptive Compression
4.1 Pruning
4.1.1 Overview of Pruning
Neural network pruning has long been recognized as an effective method for shrinking model size and speeding up model inference. The taxonomy of pruning methods can be quite complicated, including the sequential order of pruning and model training, structure specification, and the way to determine the pruned parameters [133]. Within the scope of this review, however, limiting the source model as a pre-trained large transformer for natural language processing [134, 4] or visual recognition [12, 26, 135, 136] puts forward several specific categories of techniques to be addressed (Fig. 5).
As pre-training accounts for most of the model performance for downstream tasks, pruning after training (PAT) [137] becomes a major choice. At a high level, the whole process consists of pre-training, pruning, and retraining performance recovery. However, the multi-task nature of pre-training and huge computational in the retraining phase also lead to critical issues to be addressed for pruning large transformer models, as detailed later.
On the other hand, the structure specification of pruning separates the techniques into unstructured and structured pruning. The former conducts pruning at the finest-grained level [138, 139], i.e., weight-wise pruning, which follows the optimization problem below,
$\mathop{\min}_{\mathbf{\theta}}L(\mathbf{\theta};D),s.t.\|\mathbf{\theta}\|_{0}\leq k$ (9)
where $L$ is the general loss function on the dataset $D$, $\mathbf{\theta}$ denotes the model parameter, and $k$ denotes the targeted non-zero weight number.
While the parameter size or memory usage are generally reduced in unstructured pruning, latency speed-up can not be guaranteed since the resulting model shape tends to be irregular, and acceleration requires specific hardware design. In contrast, specifying the eliminated structure as an entire layer, head, or other network unit results in structural pruning [140], which can generally shorten the latency on standard hardware. When the specification is a pattern suitable for speeding-up (e.g., a specific ratio of non-zero mask), the pruning is called semi-structural pruning [141].
Besides network parameters, reducing input size is also an appealing direction to explore. With the emphasis on removing redundant information between tokens, as both discovered in vision [33] and language domain [142] transformers, this type of algorithm also achieves decreased computation FLOPs with unpruned network parameter.
Figure 5: The taxonomy of pruning methods used for transformer models.
4.1.2 Pruning for Transformer-Based Large Language Models
Pruning granularity. In response to the considerable parameter increase in the LLM era, early attempts of pruning arise with the focus on unstructured and semi-structured pruning [143, 144, 145]. SparesGPT, for instance, conducts unstructured and semi-structured (mainly 2:4) pruning starting from OPT-175B [146] and BLOOM-176B [147] and achieve 50-60% sparsity with moderate perplexity increase and performance decrease on downstream datasets. This work demonstrates the feasibility of low-resource pruning in $>$100 B models with moderate sparsity.
Another line of work mainly focuses on the structured pruning method to achieve more significant hardware speed-up. Language structural pruning has been conducted at multiple granularity. For instance, [148] discusses the redundancy of attention head, [149] proposes a task-specific extraction of pre-trained sub-networks, [150] focuses on FFN layers, and Block Pruning [151] is introduced to prune MHA and FFN parameters separately. In comparison, CoFi [152] proposes a more comprehensive coarse- and fine-grained pruning method that learns to generate masks at different granularity, i.e., FFN layers, FFN intermediate dimensions, MHA layers, Attention heads, and hidden dimensions to encourage the optimization in high sparsity domain. In this way, the pruning of a specific parameter can be determined at different granularity in an end-to-end learnable manner. When pruning the general domain pre-trained BERT model to typical downstream datasets, CoFi achieves more than 10×10$\times$10 × speed-ups with a limited accuracy gap. Later, this method was extended to a larger decoder-based LLM with a 7B parameter size [30]. With a pre-defined model shape, the pruned model shows better instruction tuning performance than models of similar size but trained from scratch. Also, specifying the model configuration with uniform size avoids irregularities in model shape and further increases the inference throughput. In contrast to pre-defining the model shape, [29] proposed first to identify groups of coupled structures considering the interdependency of parameters in FFN, MHA, and channel-wise groups and conduct the pruning process by group. Experiments at a relatively low compression rate (20%) demonstrate the different effectiveness of these grouping strategies.
Pruning criteria. To determine the structure to be pruned, various metrics have been explored in the context of LLM pruning. For instance,  [143] follows [153] to find an optimal mask that can be used for weight reconstruction and uses Hessian Matrix (i.e., the second order indicator of loss change concerning the change of this parameter) to indicate the parameter importance, which is experimentally shown better than a magnitude criteria [154] and much faster than previous method AdaPrune [155]. In comparison, as noted by LLM Pruner [29], the first-order term should also be considered due to the shift of data used in the pruning process compared to the original language model training. Also, it proposes considering the weight importance at different levels and explores different ways of combining the importance information. Besides,  [144] also explores the combination of magnitude and norm of the input activations as the pruning criteria, which lowers the computational cost used for Hessian computation.
Learning to prune. Besides determinating parameter importance with designed metrics, one can also perform extra training, commonly with regularization on sparsity [156, 157]. In particular, the success of LLM as a multi-task language processor promotes the development of a pruning method that preserves this multi-task nature, as opposed to the task-specific trails that are more feasible in the traditional pretrain-finetune paradigm [152]. In response to this motivation, recent works such as SIMPLE [158], LLM Pruner [29] and Sheared LLama [30] incorporate the causal language modeling loss in the pruning objective. However, this “general pruning” strategy raises another issue of data selection in the pruning process, as the domain gap between pre-training and pruning may hinder the importance of estimation/performance recovery. As detailed in [30], a dynamic batch loading strategy is applied to enhance adaptively input the data batch from a proper domain to balance the model performance as a multi-task learner.
Computation cost for parameter pruning. Considering the huge parameter size, post-pruning retraining and parameter importance estimation may both generate significant computation burden for large language models. In an early attempt, [143] considers a one-shot strategy, i.e., only limited data are used to calibrate the Hessian Matrix in importance estimation, and no retraining was employed. From another perspective, LoRAPrune [159] shows that introducing parameter-efficient tuning such as Low-Rank Adaption can also decrease the computation cost in retraining while compensating for the performance degradation.
Context and token pruning. Different from the case in the computer vision domain, language sequence can be very long, and long-range reference or induction also calls for the expansion of input tokens (i.e., context) to prompt better language understanding. Also, since the inference cost of LLM is in quadratic complexity with respect to sequence length, this expansion greatly hinders the efficient inference of language models. To tackle these problems, methods have been proposed since BERT era to prune redundancy in context or context attention calculation. For context pruning,  [160, 142] proposes to prune tokens across the transformer layers progressively and improves the throughput by several folds on GLUE benchmark with less than 1% accuracy loss. Among the second line of research, various sparse attention techniques have been proposed, including local attention restricted to nearby tokens [161, 162, 163], introducing global tokens [164, 162], content-based token grouping [37, 163], and attention back tracking [165]. With the advance of the modeling capacity of LLM, the context input for LLM is also growing readily. To provide a more adaptive attention selection, [31] proposes a sparse sigmoid-based selection scheme. Also, erasing tokens from the key-value cache makes this method more hardware-friendly in a decoder-based language model. While the experiments are mainly conducted on modest size GPT-2 model [3], the empirical results that up to 80% of the context can be successfully pruned with negligible degradation in perplexity indicates an appealing potential to explore along this direction.
4.1.3 Pruning for Transformer-Based Vision Model
TABLE V: comparison of representative pruning methods on vision transformer DeiT-Base. The Speed-up ratio is calculated using the latency or throughput reported in the original paper. Top-1 Acc. is measured on ImageNet 1K. “/” indicates the keep ratio of dynamic tokens, and “-” denotes the pruning ratio of parameters.
  Models Pruning Type ##\## Params FLOPs Speed Up Top-1 Acc. (%)
DeiT-B - 86.6M 17.6G 1$\times$ 81.8%
DPS-DeiT-B [33] Unstructured (Token) 87M 9.4G 1.4$\times$ 81.5%
DynamicVIT-B/08 [166] Unstructured (Token) - 13.3G 1.2$\times$ 81.6%
[…]
SAViT-50 [168] Structured 42.6M 8.8G 1.55$\times$ 82.5%
SAViT-70 [168] Structured 25.4M 5.3G 2.05$\times$ 81.7%
ViT-Slim [32] Structured 52.6M 10.6G - 82.4%
WD-prining [169] Structured and Early-Ex 55.3M 9.9G 1.18$\times$ 80.76%
X-pruner [34] Structured 87M ${}^{\dagger}$ 8.5G - 81.02%
  • ${}^{\dagger}$ The corresponding data is from [170].
TABLE VI: comparison of representative pruning methods on large language transformer.
  Method Pruning Type Source Model Compression Speed Comparison with
Rate Up Source Model
SparseGPT [${}^{\dagger}$ Un- and semi-sturctured OPT-175B [146] 50-60% 1.54-1.79$\times$ 8.38/8.34
(WikiText2 ppl)
Sheared LLama [30] Structured LLaMA2-7B [1] 61.4% - 56.7%/64.6%
(11 datasets)
LLM Pruner [30] Structured LLaMA2-7B [1] 20% 1.18$\times$ 60.07%/68.59%
(7 datasets)
Dyna. Context Pruning[31] Context GPT-2 [3] ${}^{\star}$ 1.2$\times$ +0.085 ppl
(Wiki and bookcorpus)
  • ${}^{\dagger}$ Data is from 2:4 partial sparsity setting with GPU speeding up [143].
  • ${}^{\star}$ Compression rate is about context tokens instead of parameters.
Token and feature pruning. The early exploration of pruning for vision transformer pruning focuses on reducing information redundancy in feature space. In a preliminary study, [171] proposes learning the most informative feature dimensions and neglecting the remaining dimensions in the inference time. This is achieved by learning a diagonal real-value feature mask during pre-training with a sparsity regularizer and obtaining the hard feature mask using a defined threshold. From a different perspective, given the fact that token features in late ViT layers are quite similar in feature embedding, patch slimming [33] method is proposed to prune the patches from the output layer to the input layer. In particular, for each layer, it evaluates the impact of a patch on the final output feature and removes less important ones. Besides determining the patch importance by training dataset statistics, token sparsification can also be achieved by inserting a light-weight prediction network to indicate the importance of the input patches [33, 166, 172, 173, 167]. Note that since the self-attention can accept the token sequence regardless of its length, unstructured pruning of input tokens can be easily implemented in a hardware-friend manner for these token pruning methods. Furthermore, similar techniques are adopted in other transformer-based vision models such as DETR [174] to promote focus on object tokens [175, 176].
Structured pruning. Apart from the redundancy in feature dimension and tokens, the heavy burden of matrix computation is another important axis to improve model efficiency. In an early work, [167] proposes to selectively optimize for sparsity by obtaining unstructured subnetworks and structured pruning for attention heads, besides importance-based token selection. [169] develops a structural pruning method to reduce the network’s width by pruning linear projection matrices as well as attention heads, together with depth pruning with inserted early-exit classifiers at each layer. Later on, more works emerged to explore the collaborative effect of parameter pruning across different transformer modules. In [177], a Hessian-based criteria is proposed to evaluate the parameter importance globally over learnable parameters. Together with the parameter redistribution strategy that promotes shorter latency, this method can also introduce novel efficient ViT structures named NViT. With an emphasis on speeding up continuous structure search, [32] performs a single-shot architecture search for all the ViT components with significantly shorter GPU hours needed. Also, by incorporating the interactions between the components, [168] manages to introduce adaptively the pruning ratio for each component, adding another degree of freedom in pruning. More recently, [34] proposes to use class labels to optimize the fully differentiable pruning operation at each layer.
4.1.4 Discussion
Pruning serves as a fundamental way to reduce the computation burden of the pre-trained transformer. Traditionally, this is achieved in an orthogonal manner by pruning input features or network parameters. In particular, we notice that different pruning methods can biasedly benefit particular application purposes. For large language models, structured network pruning fits more in the setting when a modest length (e.g., several thousand tokens) is used as input to speed up inference, as the defined hardware-friendly model shape has been proven to realize a good trade-off between acceleration and perplexity decrease [30]. In comparison, context pruning or sparse attention methods deserve more attention when context length gets longer, and the ceiling of pruning ratio for context for LLMs (e.g., 7B or more) has not been determined in a detailed manner, which is an interesting direction to explore. For vision transformers, the development of pruning methods has systematically tackled each angle of model design, making the current pruning process more flexible and closer to a continuous structural search task, which can be an excellent reference to the development of language domains. The training cost is another critical axis to consider in the pruning of large models, especially when incremental LLM pre-training is considered a component of the pruning process. Therefore, developing training-efficient pruning methods is attracting more attention. And this unmet algorithm need can be fulfilled by more accurately estimating the parameter (or block) sensitivity by limited data or more detailed exploration of the information stored hidden features during training/inference.
4.2 Efficient Architecture
Transformer [11] has become the de facto backbone architecture to develop various large vision and language models (LVMs and LLMs), making it possible to scale to billions of parameters. In this section, we review the efficient architecture design of large transformer-based models in both vision and language domains, i.e., efficient design on mainstream architecture, attention mechanism and feed forward network.
TABLE VII: Model cards of several representative transformer-based LLMs and LVMs with public configuration details. Here $d_{model}$ is already expanded by expansion ratio.
  Models Arch. ##\## Params Layer Head $d_{model}$
Natural Language Processing
BERT [91] Encoder 0.1Bsimilar-$\sim$0.3B 12$\sim$24 12$\sim$16 3072$\sim$4096
PaLM [35] Decoder 9Bsimilar-$\sim$540B 32$\sim$118 16$\sim$48 4096$\sim$18432
Gopher [178] Decoder 4Msimilar-$\sim$280B 8$\sim$80 16$\sim$128 512$\sim$16384
BLOOM [147] Decoder 0.6Bsimilar-$\sim$176B 24$\sim$70 16$\sim$112 1024$\sim$14336
GPT3 [4] Decoder 175B 96 96 12288
GLM-130B [179] Decoder 130B 70 96 12288
Chinchilla [180] Decoder 70B 80 64 8192
LLaMA [1] Decoder 7Bsimilar-$\sim$65B 32$\sim$80 32$\sim$64 4096$\sim$8192
Falcon [181] Decoder 40B 60 64 8192
T5 [132] En.-De. 11B 24 128 1024
Computer Vision
ViT [127] Encoder 86Msimilar-$\sim$0.6B 12$\sim$32 12$\sim$16 3072$\sim$5120
DeiT [126] Encoder 5Msimilar-$\sim$86M 12 3$\sim$12 768$\sim$3072
TNT [182] Encoder 6Msimilar-$\sim$66M 12 3$\sim$10 768$\sim$2560
PVT [183] Pyramid 13Msimilar-$\sim$61M 12$\sim$45 8 2048
CMT [184] Pyramid 10Msimilar-$\sim$75M 24$\sim$40 8 1324$\sim$2432
Swin [38] Pyramid 29Msimilar-$\sim$0.2B 16$\sim$28 24$\sim$48 3072$\sim$6144
4.2.1 Representative Transformer-Based Models
Table VII presents the model cards of several representative LLMs and LVMs with public details. Vanilla transformer [11] is based on the encoder-decoder architecture [132]. In this architecture, the encoder stacks multi-head self-attention layers to encode the input sequence into latent representations, while the decoder performs cross-attention on these representations and generates target sequences in an autoregressive manner. Bidirectional Encoder Representations from Transformers (BERT) [91] is designed to pre-train deep bidirectional representations from unlabeled text by considering both left and right context in all layers. GPT-series models [4, 2] have effectively demonstrated the power of in-context learning using decoder-only architectures [147, 178, 181]. These decoder architectures incorporate unidirectional attention masks to ensure that each input token can only attend to previous tokens, including itself. The input and output tokens are processed in a similar manner within the decoder. In the following, we will provide separate reviews from the perspectives of CV and NLP on the current structural innovations in transformer-based architectures.
4.2.2 Language Domain
The self-attention mechanism in conventional transformer architectures often faces quadratic computational complexity. This poses a challenge for training and making inferences with long input sequences. To enhance efficiency, current innovations in the mainstream can be classified into three categories: optimizations targeting the attention mechanism, the direct replacement of attention with more efficient architectures, and enhancements focused on the Feed-Forward Network (FFN) component.
Attention Innovation. Various efficient Transformer variants have been proposed to reduce the computational complexity of the attention mechanism by incorporating structural priors on attention, such as sparsity [185, 186, 187, 188, 189, 190]. For example, Reformer [37] approximated full attention computation using locality-sensitive hashing, reducing the complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$. Other techniques include pooling-based compression [189], clustering methods [188, 191] that apply k-means clustering to learn dynamic sparse attention regions, and Longformer [192], which combines local windowed attention with task-motivated global attention. Chelba et al. [193] proposed truncating the target-side window based on an $N$-gram assumption, restricting the self-attention mechanism to use only the previous $N-1$ tokens. Additionally, locally banded sparse attention methods, such as Factorized Attention [187], have been adopted in models like GPT-3 [4].
Other methods have been proposed to reduce the computational complexity of the attention mechanism using linear attention. Performer [194] introduced fast attention via positive orthogonal random features, which is a scalable kernel method for approximating softmax attention kernels. This approach can efficiently model kernelizable attention mechanisms beyond softmax and comes with strong theoretical guarantees, including nearly unbiased estimation of the attention matrix, uniform convergence, and low estimation variance. Linear Transformer [195] reformulates self-attention as a linear dot-product of kernel feature maps. It uses kernels $\frac{\phi(q_{i})\phi(k_{j})}{\sum_{n=1}^{|x|}\phi(q_{i})\phi(k_{n})}$ to replace the softmax function and takes advantage of the associativity property of matrix products to reduce the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$.
Moreover, there have been studies [35, 196, 197] exploring the concept of multi-query attention, where different heads share the same linear transformation matrices for the keys and values, as proposed by Shazeer et al.[197]. This approach offers significant computational savings with only a minimal impact on model quality. To further strike a balance between multi-query attention and multi-head attention, GQA et al. [198] introduces to group attention heads. In this approach, heads that belong to the same group share the same transformation matrices.
FlashAttention [199] offered an innovative approach to optimizing the speed and memory utilization of attention modules on GPUs. It focuses on an Input/Output (IO)-aware perspective, making better use of the fast memory SRAM by dividing the input into blocks. This optimization technique has already been integrated into various platforms [200, 201, 202, 203, 204] developed for training LLMs.
Non-Transformer Architecture. Recent research has focused on developing new architectures for language modeling, including parameterized state space models [205, 206, 207, 208], long convolutions [209], and the incorporation of recurrent models [210, 211, 36, 212], which come with strong inductive biases. In particular, Gu et al. [205] introduced the S4 model, based on a novel parameterization for the SSM [213, 214], which involves conditioning matrix A with a low-rank correction, making it stably diagonalizable. This reduces the SSM to the well-studied computation of a Cauchy kernel [215]. The GSS model [206] further extended the use of gating units to the state space model family, and it was observed that employing gating units leads to dimensionality reduction when performing FFT operations, which addresses a major bottleneck in training speed. Hyena [209] trained a recurrence of gating units and implicitly parametrized long convolutions, which serves as an attention-free drop-in replacement for the traditional transformer architecture. Ma et al. [212] introduced MEGA, a single-head gated attention mechanism enhanced with exponential moving average. This mechanism aims to integrate the inductive bias of position-aware local dependencies into the position-agnostic attention mechanism.
Peng et al. [216] proposed RWKV based on the fact that Recurrent Neural Networks (RNNs) show linear scaling in memory and computational requirements but encounter limitations in parallelization and scalability. They leveraged a linear attention mechanism, specifically replacing the quadratic QK attention with a scalar formulation that has linear cost. Additionally, they redefined recurrence and sequential inductive biases to parallelize computations during training while maintaining constant computational and memory complexity during inference.
(a) Parallel representation.
(b) Recurrent representation.
Figure 6: Dual form of RetNet (image from [36]).
Furthermore, Retentive Network (RetNet) [36] theoretically derived the connection between recurrence and attention and proposed the retention mechanism for sequence modeling. This mechanism supports three computation paradigms, namely parallel, recurrent, and chunkwise recurrent, achieving training parallelism, low-cost inference, and good performance simultaneously. The parallel representation (also depicted in Figure 5(a)) of retention is:
$Q=(XW_{Q})\odot\Theta,$ $\quad K=(XW_{K})\odot\overline{\Theta},\quad V=XW_{V}$ (10)
$\Theta_{n}=e^{in\theta},$ $\quad D_{nm}=\left\{\begin{aligned} &\gamma^{n-m},&n\geq m\\ &0,&n<m\\ \end{aligned}\right.$
$\mathrm{Rete}$ $\mathrm{ntion}(X)=(QK^{\intercal}\odot D)V$
where $\overline{\Theta}$ is the complex conjugate of $\Theta$, and $D\in\mathbb{R}^{|x|\times|x|}$ combines causal masking and exponential decay along relative distance as one matrix. This parallel representation enables RetNet to train the models with GPUs efficiently. The recurrent version of retention (Figure 5(b)) is favorable for inference and can be written as (for $n$-th timestep):
$S_{n}=\gamma S_{n-1}+K_{n}^{\intercal}V_{n}$ (11)
$\mathrm{Rete}\mathrm{ntion}(X_{n})=Q_{n}S_{n},\quad n=1,\cdots,|x|$
where $Q,K,V,\gamma$ are the same as in Eq. 10.
FFN Innovation. Large and sparse Feed-Forward Networks (FFNs) such as Mixture-of-Experts (MoE) [217, 218, 219, 220, 221, 222] have been effective in scaling up transformer-based models for pre-training LLMs. They replace a single FFN module with multiple equally-sized modules (experts) and activate only a few experts based on the input. This selective activation of FFN improves generalization performance while maintaining fixed training and inference costs. Liu et al. [223] further found a simpler selection method known as Avg-K, which selects blocks based on their mean aggregated hidden states. This method achieved lower perplexity in language model pre-training compared to existing MoE architectures such as Switch Transformer [219] and HashLayer [220].
Romero et al. [224] presented Continuous Kernel Convolution (CKConv) for handling arbitrarily long sequences in a parallel manner within a single operation. CKConv formulates convolutional kernels as vector-valued continuous functions parameterized by a small MLP instead of a sequence of independent weights. This MLP takes a time-step as input and generates the value of the convolutional kernel at the corresponding position, allowing the generation of kernels at different resolutions and arbitrarily sizes.
PaLM [35] and LLaMA [1] both utilized the SwiGLU Activation (Swish($xW$) $\dot{x}V$) for original FFN intermediate activations. This choice was based on the observation that SwiGLU activations have been shown to substantially enhance quality in compute-equivalent experiments [225] when compared to standard activation functions such as ReLU, GeLU, or Swish [225]. It’s important to note that using SwiGLU does require three matrix multiplications in the MLP instead of two.
Architectures TP Time Memory
RNN [226] $O(Nd)$ $O(d)$
Transformer [11] $O(N^{2}d)$ $O(N^{2}+Nd)$
Linear Attention [195] $O(Nd^{2})$ $O(Nd+d^{2})$
H3 [207] $O(Nd(\log N+d))$ $O(Nd)$
S4 [205] $O(Nd^{2})$ $O(Nd)$
Hyena [209] $O(Nd(\log N+d))$ $O(N\log N\cdot d)$
Reformer [37] $O(N\log N\cdot d)$ $O(N\log N+Nd)$
Performer [194] $O(Nd^{2}\log d)$ $O((Nd+d^{2})\cdot\log d)$
RWKV [216] $O(Nd)$ $O(d)$
RetNet [36] $O(Nd^{2})$ $O(Nd)$
TABLE VIII: Model comparison from training parallelization (TP), inference cost (Time), and memory complexity (Memory). Here $N$ and $d$ denote the sequence length and feature dimension, respectively.
4.2.3 Vision Domain
Dosovitskiy et al. [127] first introduced the Vision Transformer (ViT) primarily designed for image classification tasks. In the ViT model, the input image is initially divided into a sequence of fixed-length tokens, which are then processed through multiple transformer layers to capture global relationships among the tokens. Similar to the challenges faced in the language domain, the self-attention mechanism with quadratic computational complexity is one of the major hurdles for training and inference in transformer-based architectures applied to vision tasks. Many dense prediction tasks, such as object detection and image segmentation, often require higher-resolution inputs, necessitating the manual design of efficient architectures to handle these challenges. To address these issues, novel architectures like PVT [183] and Swin [183] have been proposed. These architectures incorporate hierarchical (pyramid) structures with multiple stages, effectively overcoming the challenges of adapting the original isotropic ViT [127] to various dense prediction tasks in computer vision.
In the following, we will delve into papers that focus on improving the efficiency of ViTs, including efforts to enhance local information processing [184, 227, 228, 229, 230, 231, 232], simplify the attention mechanism [233, 234, 235, 236, 237], and explore alternative modules that can replace or work alongside attention mechanisms [238, 39, 239, 240, 241], etc.
Enhancing Locality. T2T-ViT [242] leveraged token transformations to reduce token length by iteratively aggregating neighboring tokens to one token. LeViT [243] and MobileViT [244] employed hybrid architectures with stacked convolution layers, efficiently reducing the number of features through the first layer. TNT [182] further divided original ViT’s 16$\times$16 patches into smaller 4$\times$4 patches for capturing local information. CrossViT [231] processed small-patch and large-patch tokens separately, fusing them through multiple attention mechanisms. Twins [245] alternated between local and global attention layers for improved performance. RegionViT [246] introduced regional tokens and local tokens, enhancing local context with global information. KVT [247] introduced the k-NN attention to utilize locality of images patches and ignore noisy tokens by only computing attentions with top-k similar tokens. CMT [184] applied depth-wise convolutions to augment local patterns in the attention map and intermediate activation of the FFN. Pan et al. [248] proposed HiLo attention to disentangle the high/low frequency patterns in an attention layer by separating the heads into two groups, each equipped with specialized operations that focus on local windows.
Faster Attention. Parmar et al. [233] introduced restrictions on the attention mechanism to focus on local neighborhoods. Swin [38] and Cswin [249] incorporated local attention within a window and introduced a shifted window partitioning method to enable cross-window connections. Shuffle Transformer [250] and Msg-transformer [251] employed spatial shuffle operations as alternatives to shifted window partitioning, facilitating cross-window connections. FasterViT [234] introduced hierarchical attention, breaking down global self-attention into multi-level attention components. This approach combines local window attention and hierarchical attention to achieve global information propagation while reducing computational costs. FLatten Transformer [252] integrated depth-wise convolution in conjunction with linear attention mechanisms to address the challenge of maintaining diversity in output features across different positions.
Attention-Free Architecture. AFT [253] pioneered an innovative approach by combining key and value elements with learned position biases, followed by element-wise multiplication with the query. This operation’s distinctive feature is its linear memory complexity concerning both context size and feature dimension, making it compatible with large input and model sizes. GFnet [254] presented an alternative by replacing traditional attention mechanisms with Fast Fourier Transform (FFT), frequency gating, and inverse FFT for rapid token mixing. Additionally, some architectures have been entirely based on pure multi-layer perceptrons (MLPs) [255, 40, 256, 257, 240, 241], without using convolutions or self-attention mechanisms. Yu et al. [39] proposed MetaFormer as a general architecture abstracted from ViTs without specifying the token mixer. By employing basic token mixing, primarily through non-parametric pooling, MetaFormer achieved satisfactory performance.
NAS. Indeed, the quest for optimizing transformer architectures extends to the realm of neural architecture search (NAS). Various models, including Scaling-ViT [258], ViTAS [259], AutoFormer [260] and GLiT [261], have emerged as products of NAS-driven design, demonstrating the potential for more efficient and effective transformer architectures.
4.2.4 Discussion
Transformer models consistently achieve state-of-the-art results across various tasks, but the cost of training these models, especially on long sequences, can be prohibitive. Table VIII concludes several papers that propose improvements to the attention mechanism from three different dimensions. While using linear attention can help reduce the computational complexity of the attention mechanism, it struggles to effectively encode position information, leading to a potential decrease in model performance. Models like RWKV [216] and RetNet [36], which generate outputs recursively similar to RNNs, present a promising direction for further exploration. These models only need to refer to the previous state during decoding, significantly improving decoding efficiency by avoiding the need to revisit all previous states, a characteristic of conventional Transformers. Additionally, they can encode entire sentences in parallel, enabling highly parallel and efficient training.
In contrast to NLP tasks, high-level vision tasks place a strong emphasis on the network’s ability to capture local details, contextual information, and multi-scale features for dense predictions [183]. Several studies [262, 39] have also demonstrated that the self-attention mechanism is not always indispensable for feature extraction in vision tasks. In the exploration of pure vision tasks, the possibility of directly omitting the attention module is a worthwhile avenue for further research. However, it’s worth noting that attention and cross-attention modules still play a crucial role in integrating visual features with other modalities [15]. Therefore, the quest for faster attention algorithms remains a valuable research direction.
5 Other Approaches
In addition to quantization, distillation, pruning, and novel network architectures, there are also several other model compression and acceleration approaches.
Tensor Decomposition. Tensor or matrix decomposition aims to decompose the large tensor or matrix into smaller ones so as to save the number of parameters and computational costs. This approach is firstly introduced into compression of fully-connected layers and convolutional networks [263, 264]. As for large language models, Tensor decomposition is utilized for simplifying model weight or embedding layers. Edalati et.al [265] is the first attempt to use Kronecker decomposition for compressing generative language models. This work represent weight matrix with two smaller matrices using Kronecker decomposition and achieves lossless 1.5×1.5$\times$1.5 × compression of GPT2. LoRD [266] further use low rank decomposition to compress code LLMs. TSVD [267] compress linear mappings in transformers using SVD and constrain $U$ and $V$ matrices in ternary format. In addition to model weights, TensorGPT [268] proposes to compress embedding layers and achieves more than 3×3$\times$3 × compression factor with lossless performance.
Early Exiting. Early exiting can dynamically allocate different resources for each input sample and maintain the original performance, which has been successfully used in information retrieval system [269] and convolutional networks [270]. Many early exiting technologies have been proposed for encoder-only Transformers [271, 272, 273, 274]. The key problem in early exiting is determining when to exit. The existing works mainly utilize intrinsic confidence measures [271], routing in advance [275] or training a early-exit classifier [276]. For transformers with decoders, several works have been proposed. Elbayad et.al [277] introduces the depth-adaptive transoformer for accelerating machie translation. CALM [278] proposes to dynamically allocate different computational resources for different input and timestep. CALM can achieve speedup of up to 3×3$\times$3 × on T5 encoder-decoder model. SkipDecoder [279] bypasses tokens in lower layers to middle layers which enable batch inference and reuse KV caching. SkipDecoder obtains 2×2$\times$2 × to 5×5$\times$5 × inference speedups on OPT models.
Speculative Sampling. Speculative sampling is a special accelerating approach for transformer decoding by computing several tokens in parallel [280, 281]. In large language model, decoding $K$ tokens requires $K$ runs of the model which is slow. Taking the reference tokens generated from smaller models, speculative sampling runs these tokens in parallel can significantly make the decoding process faster. Moreover, the rejection scheme [281] can preserve the distribution of the original LLM so that speculative sampling is theoretically lossless. Yang et.al [282] further take the input text as reference tokens without introducing extra models. LLMCad [283] introduces an on-device inference engine for LLMs with speculative sampling as a key technology.
Discussion. We have reviewed the methods that are different from the above in this section, including tensor decompostion, early exiting and speculative sampling. These methods do not change the basic operations, rather accelerate model inference by modifing the network connections. Tensor decompostion approximates the original weights with low-rank ones which may have performance degradation. Early exiting makes prediciton by adaptively allocate resources per sample. Speculative sampling considers the specialization of decoders and can accelerate decoding losslessly. These special methods are inspiring for developing new algorithms for accelerating LLMs or LVMs in the future.
6 Conclusions and Future directions
In this survey, we systematically investigate the compression methods for transformer models. Compared with compression method for conventional models, there are unique considerations when compressing transformer models. Unlike other architectures such as CNN or RNN, the transformer a unique architecture design with alternative attention and FFN modules, necessitating specifically tailored compression methods for optimal compression rates. Moreover, the efficiency of compression methods becomes particularly crucial for these large models. Certain model compression techniques demand substantial computational resources, which may prove prohibitive for such extensive models.This survey aims to encompass the majority of recent works pertaining to transformers and articulate a comprehensive roadmap for their compression. Subsequently, we delve into the interconnections among various methods, address post-challenges, and outline directions for future research.
Relationship between different compression methods. Different compression methods can be used together to obtain an extremely efficient architecture. A conventional sequence is to firstly define a new architecture with efficient operations. The redundant components (e.g., attention head, layers) are then removed to get a smaller model. For practical hardware implementation, quantizing weights or activations to lower bits is imperative. The choice of required bits depend on not only the tolerance of error, but also hardware design. As an illustration, Int8 computation is efficiently optimized on Nvidia A00, but it lacks support on the older Tesla P100. Distillation commonly serves as a training strategy, applicable during the fine-tuning phases of both pruning and quantization. For pursing extremely high compression rate, it is promising to explore how combining different compression strategies. Though has been well explored in conventional models like CNN [9, 284], the transformer models have more complicated architectures and higher computational cost. It is challenging to find a proper combination strategy via joint search.
Training-efficient compression strategy. In contrast to compressing conventional models, there is a heightened emphasis on the computational cost of compression methods.
Different from compression for conventional models, the computational cost for compression methods has been put on higher priority. Large transformers are currently trained on vast datasets using significant computational resources. For instance, Llama2 is trained on 2 trillion tokens with thousands of GPUs over several months [1]. It is impractical to fine-tune a model with comparable computational resources during pre-training, especially when the original data is often inaccessible. Therefore, the feasibility of efficient compression methods after training becomes more viable.
Initially developed for conventional small models, a series of works have extensively researched post-training quantization [21, 41, 22, 42], and these methods have seamlessly transitioned to transformers. With only several GPU hours, some recent works GPTQ [73] ,SmoothQuant [18] quantizes a FP16 model to Int8 without significant performance loss. However, for lower bits (e.g., 4 bit), the quantized model still surfers significant performance degradation [73, 18]. It is worth noting that extremely-low-bit models, such as binary transformers, have been extensively explored in conventional small models but remain relatively unexplored in the context of large models.
For pruning, the challenge of post-training is intricately linked to the pruning granularity. Though the nonstructural sparsity can achieve high compression rate with minimal fine-tuning requirements [143], the similar strategy is hard to transfer to the structural pruning. Directly removing the entire attention heads or layers will result in a substantial alteration of the model’s architecture and a consequent significant reduction in accuracy. How to recognize effective weights and how to effectively recover the performance are both insight directions. The efficient strategy to identify effective weights and recovery representation ability are key research directions in addressing these challenges.
Efficient architectures beyond Transformer. In real-world applications, the input context for a transformer architecture can extend to an extremely long length, encompassing sequence texts (e.g., a book with hundreds of thousands of words) in NLP or high-resolution images in CV. The vanilla attention mechanism exhibits quadratic complexity w.r.t. the length of the input sequence, posing a significant computational challenge for long-sequence inputs. Numerous studies have addressed this issue by mitigating the computational cost of attention, employing techniques such as sparse attention, local attention, etc(See Section 4.2). However, these attention compression strategies often compromise the representation ability, leading to diminished performance.
Emerging architectures such as RWKV [216] and RetNet [36] adopt a recursive output generation akin to RNNs, effectively reducing computational complexity to $\mathcal{O}(N)$. This development holds promise for further exploration in the quest for more efficient models. For computer vision tasks, even a pure MLP architecture without attention module can achieve SOTA performance [40, 241, 257, 255, 256]. Beyond the widely used Transformer architecture, it is promising to explore new efficient architecture by carefully investigating their efficiency, generalization and scaling ability.